OpenStudy (anonymous):

Write an equation for the nth term of the given geometric sequence.   –9, –27, –81, ... an i dont get how to do these types of problems

OpenStudy (jdoe0001):

well, notice that \(\bf -1^{st} \implies-9 = -3^2\\ 2^{nd} \implies-27 = -3^3\\ 3^{rd}\implies-81 = -3^4\)

OpenStudy (anonymous):

so what do you mean jdoe like you mean i should just mulitply each number or find the difference

OpenStudy (anonymous):

thats all it says jim thopm

OpenStudy (jdoe0001):

find the difference first, yes, the so-called common ratio

OpenStudy (anonymous):

a. –9(-3)^n-1 is one of the answers or selctions to choose from so idkk how to get the n-1 or what that means

OpenStudy (anonymous):

the difference is *-3 right

OpenStudy (jdoe0001):

n = nth term

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Write an equation for the nth term of the given geometric sequence. 4, 20, 100, ..., an how about this one though how do i write an equation

OpenStudy (jdoe0001):

well, the difference is just the exponent really the term 1, is 1+1 =2 for the term 2, is (1+1)+1 =3 for the term 3, is (1+1+1)+1 =4

OpenStudy (jdoe0001):

well, just find the common ratio first the geometric sequence equation is \(\bf \huge a_1 = a_1(r)^{n-1}\)

OpenStudy (jdoe0001):

\(\bf \huge a_n = a_1(r)^{n-1}\) rather for n = nth term

OpenStudy (anonymous):

is it 100(5)^n-1 or 4(5)^n-1 its gotta be one of these right? not a 4(1/5)^n-1 fractioned 1/5

OpenStudy (jdoe0001):

\(\huge a_1\) = 1st term in the sequence

OpenStudy (anonymous):

ok thank you j

OpenStudy (jdoe0001):

yw